# Kernelized Decoded Quantum Interferometry

### v2

@@ I Introduction
Quantum algorithms for discrete optimization face a fundamental dichotomy: while worst-case instances are likely intractable, practical instances often possess hidden structures—such as algebraic symmetries or sparse dependency graphs—that theoretically permit superpolynomial speedups. Decoded Quantum Interferometry (DQI) [JordanNature2025] has emerged as a powerful paradigm to exploit these structures. By encoding objective values into quantum amplitudes and subjecting them to a Fourier or Walsh–Hadamard transform, DQI attempts to concentrate the solution’s ”energy” into low-frequency modes that can be efficiently recovered by classical decoders, such as Belief Propagation or Reed-Solomon decoding. However, standard DQI relies on a fragile coincidence: the problem’s native spectral geometry must accidentally align with the decoder’s constrained capability. In the presence of hardware noise or slight structural deviations, this alignment breaks, and the ”signal” diffuses irretrievably into the high-frequency ”tail” [Bu2025Noise, Anschuetz2025RequiresStructure]. This limitation poses a critical question: Can we systematically reshape the spectrum within the quantum circuit to guarantee robustness, rather than passively hoping for favorable instance presentation?
@@
In this work, we answer affirmatively by introducing Kernelized DQI (k-DQI), a unified framework that inserts an analytically tractable, efficiently realizable unitary kernel $K$ before the interference step. Conceptually, k-DQI acts as a spectral preconditioner. Just as an optical lens focuses light to overcome transmission loss, the kernel $K$ actively remaps the objective function, concentrating its Walsh/Fourier energy onto supports that known decoders handle reliably. This architectural shift effectively converts DQI’s often implicit ”needs structure” requirement into an explicit design degree of freedom: we no longer ask binary questions like ”Does this instance have structure?”; instead, we ask ”Which kernel $K$ best exposes the structure?”
@@
The fragility of standard DQI can be phrased as a precise design question. Given a shaped spectrum $\widehat{\psi}$ and a noisy channel acting after the interference layer, the native spectral geometry must align with the capabilities of a restricted decoder family. In practice this alignment is often destroyed by hardware imperfections or slight instance perturbations. Our goal in this work is to replace this incidental alignment by an explicit, tunable degree of freedom: we seek efficiently implementable unitary kernels $K$ such that, under a realistic noise model and for a given decoder family, the resulting noise–weighted head mass $\widetilde{H}_{k}$ is provably increased and the corresponding approximation–ratio guarantees are improved. In other words, rather than asking whether an instance “has structure,” we ask which kernels expose sufficient spectral structure to place the system above a known decoding threshold.
@@
We substantiate this approach through three primary contributions that bridge quantum spectral analysis with classical coding theory. First, we establish the Monotonic Improvement Theorem. We define the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta)$, a rigorous metric quantifying the effective signal-to-noise ratio entering the decoder. We prove that under local depolarizing noise, any increase in $\Sigma_{K}$ strictly improves the k-DQI approximation ratio. This result provides a ”calculus of robustness,” ensuring that spectral engineering translates directly into algorithmic success.
@@
Second, we demonstrate structure-adaptive guarantees across distinct problem classes. For Optimal Polynomial Interpolation (OPI), we utilize chirp and Linear Canonical Transform (LCT) kernels to diagonalize algebraic symmetries, allowing k-DQI to match or exceed the performance of classical soft/list decoders. For sparse Max-XORSAT, we design block-local kernels that preserve graph locality while strictly increasing the Belief Propagation threshold (as verified by Density Evolution), effectively pushing the system back across the phase transition boundary from ”unsolvable” to ”solvable.”
@@
Finally, we show that this robustness comes at negligible computational cost. We provide explicit circuit constructions demonstrating that the kernel layer adds only linear or near-linear diagonal gates ($\tilde{O}(n)$ depth) to the pipeline. Thus, k-DQI achieves its gains purely through better spectral alignment rather than by increasing circuit depth. Collectively, these results reframe DQI from a static procedure into a tunable, noise-resilient protocol, laying the groundwork for high-dimensional generalizations governed by verifiable information-theoretic limits rather than ad-hoc heuristics.
II Kernelized DQI: model and guarantees
II.1 Pipeline and notation
@@
It coincides with the usual head mass when $\eta_{K}\equiv 1$, is nondecreasing in $d$, and is invariant under global phase. Roughly speaking, $\Sigma_{K}(\ell,\eta;d)$ measures how much of the shaped-and-kernelized spectrum survives in the head $S_{d}(\alpha)$ after passing through a noisy channel with effective per-mode attenuations $\eta_{K}(s)\in(0,1]$. A larger value of $\Sigma_{K}$ means that more probability mass reaches the decoder in a form it can reliably handle. The main message of this section is that, under a mild monotone-threshold property on the decoder, any increase in $\Sigma_{K}$ leads to a corresponding improvement in the k-DQI approximation guarantees. Theorem II.1 below makes this statement precise.
II.2 The Monotonic Improvement Theorem
@@
Theorem II.1 (Monotonic Improvement via Head Mass)
@@
Physical Intuition & Implication: Theorem II.1 transforms the algorithm design problem from an abstract search for ”quantum advantage” into a concrete optimization task: maximize the spectral energy $\Sigma_{K}$. Intuitively, the kernel $K$ acts as a ”spectral lens.” Just as an optical lens focuses light to overcome transmission loss, a well-designed kernel concentrates the problem’s solution mass into the few low-frequency modes (the ”head”) that classical decoders can reliably recover, even after attenuation by channel noise. (See Appendix A for the full operator-norm derivation and noise contraction bounds.)
ed Refer to caption
@@
II.3 Noise model and decoder property
@@
Admissible kernel families include discrete linear canonical transforms and chirp‑phase maps that (i) can be realized by poly‑depth circuits, (ii) approximately diagonalize Vandermonde‑like structure for OPI, and (iii) respect bounded‑width locality for sparse LINSAT/XORSAT so that density‑evolution assumptions hold [HealyLCT2016, Rabiner1969CZT, RichardsonUrbanke2001]. These families preserve DQI’s reduction‑to‑decoding structure and are compatible with coherent decoder blocks (e.g., reversible Gauss–Jordan/BM) [Patamawisut2025]. In IV we quantify when such kernels provably increase $\Sigma_{K}$ on OPI/LDPC‑like ensembles and how that translates to improved guarantees.
III Noise-aware bounds and limitations
@@
III.1 A noise-aware monotone bound
Under the standing conditions introduced in Sections II and II.3—polynomially bounded shaping degree, the local depolarizing–mixing–loss noise model, admissible kernels, and a decoder with a monotone threshold response—we can summarize the effect of noise on the head spectrum in a single inequality.
@@ III.1 A noise-aware monotone bound
The decoder property from Section II.3 now translates (9) directly into an approximation-ratio guarantee. There exists a nondecreasing response curve
@@
such that the k-DQI approximation-ratio lower bound $A(\ell,\eta,K;d)$ produced by the standard DQI reduction obeys
[…]
For RS-like (OPI) instances $F$ can be written explicitly using BM/KV/GS decoders [GuruswamiSudan1999, KoetterVardy2003], and for LDPC-like ensembles it is given by density-evolution or EXIT-style curves [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]. In particular, for fixed $(\ell,d,\eta,w,\mu)$, the lower bound $A(\ell,\eta,K;d)$ is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. This is the noise-aware version of the monotone theorem from Section II.2: kernels that increase the noise-weighted head mass lead to strictly stronger guarantees, up to the additive noise and coherence penalties.
III.2 Limitation: delocalized spectra
@@
IV Structure-adaptive guarantees: OPI and sparse Max-XORSAT
We now instantiate the k-DQI framework on two canonical families where decoded interferometry is particularly compelling: (i) Optimal Polynomial Interpolation (OPI) over finite fields, whose algebraic structure aligns with Reed–Solomon (RS) decoding; and (ii) sparse Max-XORSAT (a Max-LINSAT subclass) whose decoding interface is belief propagation (BP) on LDPC-like ensembles. In both cases we exhibit kernels that increase the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$ and hence improve the approximation guarantees via the monotone and noise-aware bounds from Sections II and III and the decoder response property in Section II.3.
IV.1 OPI / RS-like structure over finite fields
Let $p$ be prime and consider an OPI instance over $\mathbb{F}_{p}$ with evaluation points $\{\alpha_{i}\}_{i=1}^{m}\subset\mathbb{F}_{p}$ (all distinct) and degree bound $r$ on the unknown polynomial $h$. Following the general k-DQI pipeline, the objective encodes agreement of $h$ with constraint sets $S_{i}\subset\mathbb{F}_{p}$ and the shaped amplitude $g(x)=P(f(x))$ is defined as in (1). We take $\mathcal{F}$ to be the $p$-ary QFT and choose kernels from discrete linear canonical transform (LCT) families, i.e., unitaries generated by quadratic-phase multiplications and fractional Fourier maps (see, e.g., [HealyLCT2016, Rabiner1969CZT]). These kernels admit efficient circuit realizations and remain compatible with coherent decoding modules [Patamawisut2025].
@@ IV.1 OPI / RS-like structure over finite fields
for some parameter choice $\theta^{\star}$. Under PPC, local noise and mixing reduce the head mass according to the noise model of Section II.3 and the contraction bounds in Appendix B, but the weighted head mass remains large as long as $d=d_{\star}$ and $p$ is moderate. The following theorem packages this into a k-DQI guarantee.
OPI case: kernelized RS soft decoding.
In the OPI setting we can now combine the PPC property with the noise-aware bounds of Section III to obtain an explicit RS-type guarantee for k-DQI. Assume the PPC condition of Appendix C for some degree bound $r$ and a head size $d_{\star}=O(r)$, and consider the local depolarizing–mixing–loss noise model introduced in Section II.3. Then for head size $d=d_{\star}$ there exists a chirp/LCT kernel $K(\theta^{\star})$ such that the noise-weighted head-spectrum mass after interference satisfies
@@
where $F_{\mathrm{RS}}$ is the response function induced by standard algebraic soft/list decoders (BM/KV/GS) for RS-like families [GuruswamiSudan1999, KoetterVardy2003, ReedSolomon1960]. In particular, as soon as the effective head mass $(1-\varepsilon_{r,p})\,\underline{\eta}-\Delta(w)$ exceeds the KV/BM decoding threshold, the k-DQI pipeline with kernel $K(\theta^{\star})$ achieves a strictly better approximation-ratio lower bound than unkernelized DQI with $K=I$; this comparison is with respect to the monotone and noise-aware DQI bounds developed in Sections II and III, and follows directly from the monotonicity of these bounds in $\Sigma_{K}$.
Figure 3 empirically illustrates this behaviour for a typical quadratic OPI instance. As the chirp parameter $\theta$ is scanned, the noise-weighted head mass $\Sigma_{K}(\ell,\eta;d_{\star})$ forms a pronounced peak around an optimal value $\theta_{\star}$, where the quadratic component of the kernel nearly cancels the polynomial phase and concentrates almost all spectral weight on the head modes. Only this narrow window of $\theta$ lifts $\Sigma_{K}$ above the KV/BM soft-decoding threshold (dashed line), whereas for most kernel choices the head mass remains far below threshold. Together with the monotone theorem in $\Sigma_{K}$ and the noise-aware bound (Theorems II.1 and B.1), this shows that tuning a single kernel parameter is sufficient to move k-DQI between non-advantage and advantage regimes on OPI instances.
Figure 3: Noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d_{\star})$ for a representative OPI instance as a function of the chirp/LCT kernel parameter $\theta$. The curve exhibits a sharp peak around an optimal parameter $\theta_{\star}$, where the quadratic kernel cancels the OPI polynomial phase and concentrates almost all spectral mass in the head. The horizontal dashed line marks a KV/BM soft-decoding threshold: only a narrow neighbourhood of $\theta_{\star}$ pushes $\Sigma_{K}$ above this level, illustrating the tunable head-mass boost asserted in Theorem IV.1.
@@
IV.2 Sparse Max-XORSAT / LDPC-like structure
@@
Before stating a structural guarantee, we validate the DE surrogate by finite-length BP scans on BSC and AWGN channels. Figure 4 shows the characteristic S-shaped FER curves for a $(3,6)$ ensemble and marks the corresponding BP thresholds in erasure/noise level. These thresholds are the numerical counterparts of the decoder response curves $F_{\mathrm{LDPC}}$ introduced in Section II.3: they indicate where the density-evolution fixed point drops below the diagonal and hence connect a local head-mass lift $\Delta\Sigma_{\mathrm{loc}}$ to observable BP performance.
@@
LDPC case: block-local kernels shift BP thresholds.
In the LDPC-like setting we work under the standing assumptions of Sections II and III together with the block-local alignment (BLA) property of Appendix D. We choose the head size $d=\Theta(n/b)$ so that the k-DQI head budget assigns a constant number of dominant modes per block. For the local depolarizing–mixing–loss noise model of Section II.3, BLA guarantees that replacing the identity kernel $K=I$ by a block-local kernel $K$ increases the pre-noise head mass by an additive amount $\Delta\Sigma_{\mathrm{loc}}$. Combining this with the per-mode contraction and leakage analysis of Appendix B yields a corresponding increase in the noise-weighted head-spectrum mass,
@@
Unlike the OPI setting, LDPC-like ensembles do not admit global diagonalization by simple kernels. However, block-local alignment suffices to add head mass in a controlled, extensive way while preserving sparsity assumptions. The analysis above IV.2, together with the finite-length and cross-ensemble thresholds in Figures 4 and 5, shows that such block-local kernels manifest as rightward shifts of DE/EXIT thresholds proportional to $\Delta\Sigma_{\mathrm{loc}}$, which, via the noise-aware lower bound, translates into improved k-DQI guarantees.
V Circuit cost and efficiency

### v3

@@ 1 Introduction
Quantum algorithms for discrete optimization face a fundamental dichotomy: while worst-case instances are likely intractable, practical instances often possess hidden structures—such as algebraic symmetries or sparse dependency graphs—that theoretically permit superpolynomial speedups. Decoded Quantum Interferometry (DQI) [JordanNature2025] has emerged as a powerful paradigm to exploit these structures. By encoding objective values into quantum amplitudes and subjecting them to a Fourier or Walsh–Hadamard transform [NielsenChuang2010, Coppersmith2002AQFT], DQI attempts to concentrate the solution’s “energy” into low-frequency modes that can be efficiently recovered by classical decoders, such as belief propagation or Reed–Solomon decoding [Kschischang2001Factor, RichardsonUrbanke2001, Gallager1962LDPC, MacKay1999Good, Sudan1997RS, GuruswamiSudan1999, Koetter2003ASD]. However, standard DQI relies on a fragile coincidence: the problem’s native spectral geometry must accidentally align with the decoder’s constrained capability. In the presence of hardware noise or slight structural deviations, this alignment breaks, and the “signal” diffuses irretrievably into the high-frequency tail [Anschuetz2025RequiresStructure, Preskill2018NISQ, McClean2018Barren, Wang2021NIBP]. This limitation poses a critical question: Can we systematically reshape the spectrum within the quantum circuit to guarantee robustness, rather than passively hoping for favorable instance presentation?
@@
In this work, we answer affirmatively by introducing Kernelized DQI (k-DQI), a unified framework that inserts an analytically tractable, efficiently realizable unitary kernel $K$ before the interference step. Conceptually, k-DQI acts as a spectral preconditioner. Just as an optical lens focuses light to overcome transmission loss, the kernel $K$ actively remaps the objective function, concentrating its Walsh/Fourier energy onto supports that known decoders handle reliably [Kschischang2001Factor, RichardsonUrbanke2001, Gallager1962LDPC, MacKay1999Good]. This architectural shift effectively converts DQI’s often implicit “needs structure” requirement into an explicit design degree of freedom: we no longer ask binary questions like “Does this instance have structure?”; instead, we ask “Which kernel $K$ best exposes the structure?”
@@
We will refer to this use of $K$ as spectral preconditioning. It is helpful to view our setting as a reusable design template: (i) choose a shaping map $P(f)$ that embeds the objective into a spectrum compatible with a fixed interferometer $F$; (ii) insert a parameterized kernel family $K(\theta)$ in front of $F$; and (iii) tune $\theta$ so as to maximize a simple spectral statistic such as the noise–weighted head mass $\Sigma_{K}(\ell,\eta;d)$. Therefore, the fragility of standard DQI can be phrased as a precise design question. Given a shaped spectrum $\widehat{\psi}$ and a noisy channel acting after the interference layer, the native spectral geometry must align with the capabilities of a restricted decoder family. In practice this alignment is often destroyed by hardware imperfections or slight instance perturbations [Preskill2018NISQ]. Our goal in this work is to replace this incidental alignment by an explicit, tunable degree of freedom: we seek efficiently implementable unitary kernels $K$ such that, under a realistic noise model and for a given decoder family, the resulting noise-weighted head mass $\widetilde{H}_{k}$ is provably increased and the corresponding approximation-ratio guarantees are improved. In other words, rather than asking whether an instance “has structure,” we ask which kernels expose sufficient spectral structure to place the system above a known decoding threshold [RichardsonUrbanke2001].
@@
We substantiate this approach through three primary contributions that bridge quantum spectral analysis with classical coding theory. First, we establish the Monotonic Improvement Theorem. We define the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta)$, a rigorous metric quantifying the effective signal-to-noise ratio entering the decoder, and prove that under local depolarizing noise [Preskill2018NISQ] any increase in $\Sigma_{K}$ strictly improves the k-DQI approximation ratio. This result provides a “calculus of robustness,” ensuring that spectral engineering translates directly into algorithmic success.
@@
Second, we demonstrate structure-adaptive guarantees across distinct problem classes. For Optimal Polynomial Interpolation (OPI), we utilize chirp and Linear Canonical Transform (LCT) kernels to diagonalize algebraic symmetries [Namias1980FRFT, Healy2016LCTBook, Rabiner1969CZT], allowing k-DQI to match or exceed the performance of classical soft/list decoders [Sudan1997RS, GuruswamiSudan1999, Koetter2003ASD]. For sparse Max-XORSAT, we design block-local kernels that preserve graph locality while strictly increasing the belief-propagation threshold (as verified by density evolution) [Kschischang2001Factor, RichardsonUrbanke2001], effectively pushing the system back across the phase-transition boundary from “unsolvable” to “solvable.”
@@
Finally, we show that this robustness comes at negligible computational cost. We provide explicit circuit constructions demonstrating that the kernel layer adds only linear or near-linear diagonal/quadratic-phase gates to the pipeline [NielsenChuang2010, Coppersmith2002AQFT]. Thus, k-DQI achieves its gains purely through better spectral alignment rather than by increasing circuit depth. Collectively, these results reframe DQI from a static procedure into a tunable, noise-resilient protocol, laying the groundwork for high-dimensional generalizations governed by verifiable information-theoretic limits rather than ad-hoc heuristics.
2 Results
2.1 Kernelized DQI: model and guarantees
@@
It coincides with the usual head mass when $\eta_{K}\equiv 1$, is nondecreasing in $d$, and is invariant under global phase. Roughly speaking, $\Sigma_{K}(\ell,\eta;d)$ measures how much of the shaped-and-kernelized spectrum survives in the head $S_{d}(\alpha)$ after passing through a noisy channel with effective per-mode attenuations $\eta_{K}(s)\in(0,1]$. A larger value of $\Sigma_{K}$ means that more probability mass reaches the decoder in a form it can reliably handle. The main message of this section is that, under a mild monotone-threshold property on the decoder, any increase in $\Sigma_{K}$ leads to a corresponding improvement in the k-DQI approximation guarantees. Theorem 1 below makes this statement precise.
2.2 The Monotonic Improvement Theorem
@@
Theorem 1 (Monotonic Improvement via Head Mass).
@@
Physical Intuition & Implication: Theorem 1 transforms the algorithm design problem from an abstract search for ”quantum advantage” into a concrete optimization task: maximize the spectral energy $\Sigma_{K}$. Intuitively, the kernel $K$ acts as a ”spectral lens.” Just as an optical lens focuses light to overcome transmission loss, a well-designed kernel concentrates the problem’s solution mass into the few low-frequency modes (the ”head”) that classical decoders can reliably recover, even after attenuation by channel noise. (See Appendix A for the full operator-norm derivation and noise contraction bounds.)
@@
2.3 Noise model and decoder property
@@
Admissible kernel families include discrete linear canonical transforms and chirp‑phase maps that (i) can be realized by poly‑depth circuits, (ii) approximately diagonalize Vandermonde‑like structure for OPI, and (iii) respect bounded‑width locality for sparse LINSAT/XORSAT so that density‑evolution assumptions hold [Healy2016LCTBook, Rabiner1969CZT, RichardsonUrbanke2001]. These families preserve DQI’s reduction‑to‑decoding structure and are compatible with coherent decoder blocks (e.g., reversible Gauss–Jordan/BM) [Patamawisut2025]. In 2.5 we quantify when such kernels provably increase $\Sigma_{K}$ on OPI/LDPC‑like ensembles and how that translates to improved guarantees.
At a more abstract level, the k-DQI architecture can be viewed as an instance of a broader “spectral preconditioning” template that no longer depends on the details of decoded interferometry. The template consists of three conceptual modules:
  1. 1.
[…]
  2. 2.
    Kernelized evolution. Insert a unitary kernel family $\{K(\theta)\}_{\theta\in\Theta}$ in front of $F$. The family should (i) be efficiently implementable, (ii) respect locality constraints such as factor-graph structure on sparse instances, and (iii) contain the identity $K=\mathbb{I}$ as a baseline.
  3. 3.
    Spectral metric and tuning. Evaluate a coarse-grained spectral statistic that summarizes how much decoder-relevant probability reaches the “head” modes after the noisy channel. In this work the central quantity is the noise-weighted head mass $\Sigma_{K}(\ell,\eta;d)$ in Eq. (4). The monotonic improvement theorem (Theorem II.1) shows that the rigorous approximation-ratio lower bound $A(\ell,\eta,K;d)$ is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$, so maximizing this scalar directly improves the certified performance.
Viewed in this way, k-DQI instantiates a general design rule: rather than trying to engineer the full wavefunction, one couples the circuit to a low-dimensional spectral objective that is provably aligned with the downstream decoder. Nothing in the recipe above is inherently tied to interferometry. In QAOA-style algorithms, the shaping step is implemented by the choice of cost and mixer Hamiltonians and their angles, while a kernel layer could be used to concentrate spectral weight into subspaces where classical rounding performs best. Likewise, quantum singular value transformation and other Fourier-based algorithms already implement polynomial filters of block-encoded operators; here, an additional preconditioning layer optimized against a proxy such as $\Sigma_{K}$ or an analogous filter norm could trade increased robustness for modest circuit overhead. In the present work we make this design principle concrete for decoded interferometry, but the same template is naturally compatible with any setting where performance is governed by spectral concentration and threshold phenomena.
@@
2.4 A noise-aware monotone bound
Under the standing conditions introduced in Sections 2 and 2.3—polynomially bounded shaping degree, the local depolarizing–mixing–loss noise model, admissible kernels, and a decoder with a monotone threshold response—we can summarize the effect of noise on the head spectrum in a single inequality.
@@ 2.4 A noise-aware monotone bound
The decoder property from Section 2.3 now translates (9) directly into an approximation-ratio guarantee. There exists a nondecreasing response curve $F(\,\cdot\,)$ determined by the decoder family and problem ensemble such that the k-DQI approximation-ratio lower bound $A(\ell,\eta,K;d)$ produced by the standard DQI reduction obeys
@@
For RS-like (OPI) instances $F$ can be written explicitly using BM/KV/GS decoders [GuruswamiSudan1999, KoetterVardy2003], and for LDPC-like ensembles it is given by density-evolution or EXIT-style curves [RichardsonUrbanke2001, RichardsonShokrollahiUrbanke2001, Kschischang2001Factor]. In particular, for fixed $(\ell,d,\eta,w,\mu)$, the lower bound $A(\ell,\eta,K;d)$ is a nondecreasing function of $\Sigma_{K}(\ell,\eta;d)$. This is the noise-aware version of the monotone theorem from Section 2.2: kernels that increase the noise-weighted head mass lead to strictly stronger guarantees, up to the additive noise and coherence penalties.
@@
2.5 Structure-adaptive guarantees: OPI and sparse Max-XORSAT
We now instantiate the k-DQI framework on two canonical families where decoded interferometry is particularly compelling: (i) Optimal Polynomial Interpolation (OPI) over finite fields, whose algebraic structure aligns with Reed–Solomon (RS) decoding; and (ii) sparse Max-XORSAT (a Max-LINSAT subclass) whose decoding interface is belief propagation (BP) on LDPC-like ensembles. In both cases we exhibit kernels that increase the noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d)$ and hence improve the approximation guarantees via the monotone and noise-aware bounds from Sections 2 and 2.4 and the decoder response property in Section 2.3.
OPI / RS-like structure over finite fields
Let $p$ be prime and consider an OPI instance over $\mathbb{F}_{p}$ with evaluation points $\{\alpha_{i}\}_{i=1}^{m}\subset\mathbb{F}_{p}$ (all distinct) and degree bound $r$ on the unknown polynomial $h$. Following the general k-DQI pipeline, the objective encodes agreement of $h$ with constraint sets $S_{i}\subset\mathbb{F}_{p}$ and the shaped amplitude $g(x)=P(f(x))$ is defined as in (1). We take $\mathcal{F}$ to be the $p$-ary QFT and choose kernels from discrete linear canonical transform (LCT) families, i.e., unitaries generated by quadratic-phase multiplications and fractional Fourier maps (see, e.g., [Healy2016LCTBook, Rabiner1969CZT]). These kernels admit efficient circuit realizations and remain compatible with coherent decoding modules [Patamawisut2025].
@@ 2.5 Structure-adaptive guarantees: OPI and sparse Max-XORSAT
for some parameter choice $\theta^{\star}$. Under PPC, local noise and mixing reduce the head mass according to the noise model of Section 2.3 and the contraction bounds in Appendix B, but the weighted head mass remains large as long as $d=d_{\star}$ and $p$ is moderate. The following theorem packages this into a k-DQI guarantee.
In the OPI setting we can now combine the PPC property with the noise-aware bounds of Section 2.4 to obtain an explicit RS-type guarantee for k-DQI. Assume the PPC condition of Appendix C for some degree bound $r$ and a head size $d_{\star}=O(r)$, and consider the local depolarizing–mixing–loss noise model introduced in Section 2.3. Then for head size $d=d_{\star}$ there exists a chirp/LCT kernel $K(\theta^{\star})$ such that the noise-weighted head-spectrum mass after interference satisfies
@@
where $F_{\mathrm{RS}}$ is the response function induced by standard algebraic soft/list decoders (BM/KV/GS) for RS-like families [GuruswamiSudan1999, KoetterVardy2003, ReedSolomon1960]. In particular, as soon as the effective head mass $(1-\varepsilon_{r,p})\,\underline{\eta}-\Delta(w)$ exceeds the KV/BM decoding threshold, the k-DQI pipeline with kernel $K(\theta^{\star})$ achieves a strictly better approximation-ratio lower bound than unkernelized DQI with $K=I$; this comparison is with respect to the monotone and noise-aware DQI bounds developed in Sections 2 and 2.4, and follows directly from the monotonicity of these bounds in $\Sigma_{K}$.
Figure 3 empirically illustrates this behaviour for a typical quadratic OPI instance. As the chirp parameter $\theta$ is scanned, the noise-weighted head mass $\Sigma_{K}(\ell,\eta;d_{\star})$ forms a pronounced peak around an optimal value $\theta_{\star}$, where the quadratic component of the kernel nearly cancels the polynomial phase and concentrates almost all spectral weight on the head modes. Only this narrow window of $\theta$ lifts $\Sigma_{K}$ above the KV/BM soft-decoding threshold (dashed line), whereas for most kernel choices the head mass remains far below threshold. Together with the monotone theorem in $\Sigma_{K}$ and the noise-aware bound (Theorems 1 and 4), this shows that tuning a single kernel parameter is sufficient to move k-DQI between non-advantage and advantage regimes on OPI instances.
Figure 3: Noise-weighted head-spectrum mass $\Sigma_{K}(\ell,\eta;d_{\star})$ for a representative OPI instance as a function of the chirp/LCT kernel parameter $\theta$. The curve exhibits a sharp peak around an optimal parameter $\theta_{\star}$, where the quadratic kernel cancels the OPI polynomial phase and concentrates almost all spectral mass in the head. The horizontal dashed line marks a KV/BM soft-decoding threshold: only a narrow neighbourhood of $\theta_{\star}$ pushes $\Sigma_{K}$ above this level, illustrating the tunable head-mass boost
@@
Sparse Max-XORSAT / LDPC-like structure
@@
Before stating a structural guarantee, we validate the DE surrogate by finite-length BP scans on BSC and AWGN channels. Figure 4 shows the characteristic S-shaped FER curves for a $(3,6)$ ensemble and marks the corresponding BP thresholds in erasure/noise level. These thresholds are the numerical counterparts of the decoder response curves $F_{\mathrm{LDPC}}$ introduced in Section 2.3: they indicate where the density-evolution fixed point drops below the diagonal and hence connect a local head-mass lift $\Delta\Sigma_{\mathrm{loc}}$ to observable BP performance.
@@
In the LDPC-like setting we work under the standing assumptions of Sections 2 and 2.4 together with the block-local alignment (BLA) property of Appendix D. We choose the head size $d=\Theta(n/b)$ so that the k-DQI head budget assigns a constant number of dominant modes per block. For the local depolarizing–mixing–loss noise model of Section 2.3, BLA guarantees that replacing the identity kernel $K=I$ by a block-local kernel $K$ increases the pre-noise head mass by an additive amount $\Delta\Sigma_{\mathrm{loc}}$. Combining this with the per-mode contraction and leakage analysis of Appendix B yields a corresponding increase in the noise-weighted head-spectrum mass,
@@
Unlike the OPI setting, LDPC-like ensembles do not admit global diagonalization by simple kernels. However, block-local alignment suffices to add head mass in a controlled, extensive way while preserving sparsity assumptions. The analysis above 2.5, together with the finite-length and cross-ensemble thresholds in Figures 4 and 5, shows that such block-local kernels manifest as rightward shifts of DE/EXIT thresholds proportional to $\Delta\Sigma_{\mathrm{loc}}$, which, via the noise-aware lower bound, translates into improved k-DQI guarantees.
[…]
On this unstructured ensemble the bare DQI pipeline (identity kernel, no problem-specific decoder) actually performs slightly worse than the Monte–Carlo baseline with the same shot budget. This behavior is fully consistent with Theorem 1 and the noise-aware lower bound in Theorem 4: when the shaped spectrum is effectively delocalized, the identity kernel produces a very small head mass $\Sigma_{K}$ and there is no reason for $A(\ell,\eta,K;d)$ to exceed the classical baseline. It also aligns with recent “no-advantage” observations for DQI on random CSP families and MaxCut [Anschuetz2025RequiresStructure, Parekh2025MaxCut]. In contrast, once a simple chirp kernel is tuned to increase $\Sigma_{K}$, the k-DQI curve in Figure 6 stays uniformly above both the Monte–Carlo and bare-DQI baselines, illustrating how spectral preconditioning can convert an abstract head-mass gain into a concrete finite-size performance boost even on small random instances.
Figure 6: Spectral preconditioning on an unstructured Max-XORSAT ensemble. We consider random Max-XORSAT instances with $n=10$ variables and $m=20$ clauses and compare three strategies under the same measurement budget of $M=200$ shots per instance: a classical Monte–Carlo baseline that samples bit strings uniformly at random, standard DQI with the identity kernel, and k-DQI with a global chirp kernel $K(\gamma)$. For k-DQI we scan a one-dimensional grid of chirp parameters $\gamma$ and retain the kernel that maximizes the pre-noise noise-weighted head mass $\Sigma_{K}(\ell,\eta;d)$ at shaping degree $\ell=2$. For each value of the mixing parameter $\epsilon$ we plot the average approximation ratio $\langle f\rangle/f_{\mathrm{opt}}$ over $30$ random instances, where $f_{\mathrm{opt}}$ denotes the true optimum of that instance; error bars show the standard error across instances.
3 Methods